# ZTF SN Ia DR2: Overview

### v1

@@
We present the first homogeneous release of several thousand Type Ia supernovae (SNe Ia), all having spectroscopic classification, and spectroscopic redshifts for half the sample. This release, named the “DR2”, contains 3628 nearby () SNe Ia discovered, followed and classified by the Zwicky Transient Facility survey between March 2018 and December 2020. Of these, 3000 have good-to-excellent sampling and 2667 pass standard cosmology light-curve quality cuts. This release is thus the largest SN Ia release to date, increasing by an order of magnitude the number of well characterized low-redshift objects. With the “DR2”, we also provide a volume-limited () sample of nearly a thousand SNe Ia. With such a large, homogeneous and well controlled dataset, we are studying key current questions on SN cosmology, such as the linearity SNe Ia standardization, the SN and host dependencies, the diversity of the SN Ia population, and the accuracy of the current light-curve modeling. These, and more, are studied in detail in a series of articles associated with this release. Alongside the SN Ia parameters, we publish our force-photometry gri-band light curves, 5138 spectra, local and global host properties, observing logs, and a python tool to ease use and access of these data. The photometric accuracy of the “DR2” is not yet suited for cosmological parameter inference, which will follow as “DR2.5” release. We nonetheless demonstrate that the multi-thousand SN Ia Hubble Diagram has a typical 0.15 mag scatter.
@@
For the last thirty years, Type Ia Supernovae (SNe Ia) have played a central role in building the current standard model of cosmology. In the late 90s, SNe Ia led to the discovery of the acceleration of the expansion rate of the Universe, the cause of which was dubbed dark energy (Riess et al., 1998; Perlmutter et al., 1999), see Goobar & Leibundgut (2011) for a review. The existence of dark energy has since been confirmed with high precision by various other cosmological probes, such as the anisotropies in the cosmic microwave background (Spergel et al., 2003; Planck Collab., 2020) and from the baryon acoustic oscillations (e.g., Eisenstein et al., 2005; BOSS Collab., 2017). The two decades that followed were those of the SN cosmology field maturation. Many low- () and high-redshift () surveys have enabled to gather, together, SNe Ia. Joining these datasets, the dark energy equation of state parameter has been shown to be in good agreement with , expected if the acceleration of the Universe’s expansion is due to a simple cosmological constant in Einstein’s general relativity equations (Astier et al., 2006; Betoule et al., 2014; Scolnic et al., 2018; Brout et al., 2022). This gain in statistics allows us to test the use of SNe Ia as accurate standard candles of cosmology, since, despite the great success of SN cosmology, their underlying astrophysics are still largely unknown.
@@
It is generally accepted that a SN Ia is the transient event resulting from the thermonuclear-explosion of a carbon-oxygen white dwarf (Whelan & Iben, 1973; Iben & Tutukov, 1984; Nugent et al., 2011), triggered by accreting material from a companion star in a binary system (Liu et al., 2023). However, the nature of the companion star (another white dwarf, main-sequence star, etc.) and the explosion mechanism (at or below the Chandrasekhar mass limit) is still unclear and no single picture has emerged (Maoz et al., 2014). In recent years, sample studies of SNe Ia have been used, as they provide a population-wide approach to probe various aspects of SNe Ia physics (e.g. Maguire et al., 2014; Silverman et al., 2015; Papadogiannakis et al., 2019; Tucker et al., 2020; Desai et al., 2023). However, these analyses are so far limited by small-number statistics and/or survey design (e.g., coverage, cadence, depth). The astrophysical origin of SNe Ia, and their homogeneity, are consequently still highly debated (Jha et al., 2019). Additional data, notably spectroscopic and close to the explosion epoch, would be valuable to discriminate competitive models (e.g., Deckers et al., 2022, and references there in). Nonetheless, understanding the SN Ia mechanism is not an actual requirement for SN cosmology, as long as the astrophysical dependencies can be controlled below the statistical uncertainties. But, without a better grasp on the actual SN Ia astrophysics, such an assertion is difficult.
@@
Since the mid-90s, the success of SN cosmology emerged from our ability to standardize their brightness from a natural scatter of down to by exploiting two empirical linear relations that correlate the SN Ia’s stretch and color, derived from their light-curve, to their absolute peak brightness (Riess et al., 1996; Tripp, 1998; Guy et al., 2010). A scatter, i.e., 7% precision in distance, makes it one of the most precise standard candles in astrophysics. However, only half of this variance can be explained by measurement or light curve modeling errors. Thus, to further control any unexplained astrophysical dependency in SN Ia distances, the correlation between the SN properties and that of their host environments have been extensively studied over the past decade.
@@
Early on, it has been suggested that two populations of SNe Ia must coexist to explain the relative rates between progenitors (e.g., Mannucci et al., 2005; Sullivan et al., 2006; Smith et al., 2012): one population follows recent star formation (called ‘prompt‘) with the other related to stellar mass (i.e, old population stars, called ‘delayed‘). More delicate for cosmology, the stretch- and color-standardized SN Ia magnitude has then been shown to depend on its environment, such that those from low-mass hosts are fainter then those from high-mass galaxies (e.g., Sullivan et al., 2010; Childress et al., 2013; Roman et al., 2018). This so called mass-step has since been used as a third standardization parameters in all cosmological analysis (e.g., Betoule et al., 2014; Scolnic et al., 2018; Brout et al., 2022; Riess et al., 2022). Yet, the origin of this bias is still highly debated. An inaccurate correction of this effect may lead to significant bias in the derivation of cosmological parameters (Rigault et al., 2015, 2020; Smith et al., 2020). Currently, the most discussed models are: differences between host interstellar dust properties (Brout & Scolnic, 2021; Popovic et al., 2023) or progenitor age, in the context of the prompt vs. delayed model (Rigault et al., 2013, 2020; Nicolas et al., 2021; Briday et al., 2022). Both may actually be true (e.g., Kelsey et al., 2021; Wiseman et al., 2022).
@@
Today, state-of-the-art compilations of samples nearly reach 2000 SNe Ia and are limited by systematic uncertainties (Brout et al., 2022; Vincenzi et al., 2024). Along side the astrophysical biases, the dominating source of known uncertainties are calibration issues largely due to the requirement to compile SN Ia samples from multiple surveys spanning various redshift ranges (e.g., Betoule et al., 2014; Brout et al., 2022; Vincenzi et al., 2024). This issue is particularly critical at low-redshift where dozen of samples are merged, the largest of which contains less than 200 targets (e.g., Brout et al., 2022).
@@
After two decades, the standard model of cosmological starts to see first hints of inconsistencies. and, again, SN Ia play a key role. First and foremost, the direct measurement of the Hubble constant derived by anchoring the absolute SN Ia standardized luminosity with Cepheids is found to be higher than theoretical expectations anchored by cosmic microwave background data (e.g., Riess et al., 2016, 2022; Planck Collab., 2020). Most recently, two independent SN Ia compilations, concluded that might differ from at the 2 to level (Rubin et al., 2023; DES Collab., 2024). These results, strengthen by the recent DESI Year-1 release (DESI Collab., 2024), could be the sign of new fundamental physics, or hints of unknown systematic biases in the SN Ia distances.
@@
In that context, the Zwicky Transient Facility (ZTF, Bellm et al., 2019; Graham et al., 2019) has been collecting thousands of nearby SNe a year in the Northern sky since it started science operation in March 2018. As of mid-2024 ZTF-discovered SNe have been spectroscopically classified. In this paper we present an overview of the second data-release associated to the ”Type Ia Supernovae & Cosmology” science working group, aka “DR2”. This dataset contains 3628 Type Ia Supernovae (SNe Ia) detected, followed and classified by the ZTF survey and discovered before the end of December 2020.
@@
This “DR2” follows Dhawan et al. (2022) where we illustrated the characteristics of the survey using the first months of operation. Along with this release overview, 20 companion papers that study the dataset, providing key insights on nearby SN Ia physics and their use as cosmological probes are released. However, we warn the user that the current light curve photometry calibration does not yet reach the accuracy to unlock cosmological analysis. Any other scientific study is welcomed. Ongoing work is been finished on the survey calibration that will lead to a cosmological parameter inference as a next release: “DR2.5”.
@@
With 2667 well sampled spectroscopic nearby SNe Ia passing usual light curve quality cuts, this release is the largest SN Ia release to date, over any redshift range. It increases by an order of magnitude the current state-of-the-art low-redshift sample compilation of 192 targets used in DES Collab. (2024) to anchor their Hubble diagram, and by five the number of SN Ia ever used for cosmology (Scolnic et al., 2022).
@@
We summarize in Section 2 the ZTF survey operation during the period covered by this release. In Section 3, we introduce the “DR2” sample and we briefly review the release data in Section 4. We review in Section 5 of the 20 companion papers, while the content and access of the release is summarized in Section 6. We conclude in Section 7.
@@ 2 Summary of the ZTF operation
The ZTF survey employs the ZTF-camera mounted on the P48 Schmidt telescope at the Mount Palomar Observatory. As detailed in Dekany et al. (2020), this 576 megapixel camera is made of sixteen 6144×6160 e2v CCD231-C6 CCDs and equipped with three filters: ztf:g (), ztf:r () and ztf:i (). Each CCD is subdivided into four read-out channels (called quadrants) that have square 1.01 arcsec pixels, selected to match the typical site image quality of 2 arcsec full-width at half-maximum (FWHM; see Smith et al. ). The camera has a 47 deg2 field-of-view with a 86.7% filling factor. It reaches a typical 20.5 mag -limit depth in 30 s, and, with a slightly less than 9 s read-out overhead made while slewing, ZTF is able to cover, at that depth, 3750 deg2/hour. Between March 2018 and December 2020, ZTF has acquired 480,572 exposures, with a typical same-filter cadence of three days, thanks to the National Science Foundation (NSF) funded Mid-Scale Innovations Program (MSIP) public survey program (Bellm et al., 2019). During the DR2 operations, the ZTF collaboration also operated a number of additional surveys, including an extragalactic high-cadence survey, with six visits a night in the g and r bands, a larger area ( deg2) survey that acquired same-night exposures, mostly in and , and an i-band survey with a four-day cadence (Bellm et al., 2019).
ZTF also has a low-resolution integral field spectrograph (SEDm, Blagorodnova et al., 2018; Rigault et al., 2019; Kim et al., 2022; Lezmy et al., 2022) dedicated to spectroscopically classify transients detected by the photometric survey. The Bright Transient Survey (BTS) ZTF cross working group is designed to use the SEDm, and other spectrographs, to construct a magnitude-limited spectroscopically classified SN sample. As detailed in Perley et al. (2020), they reach a 97%, 93% and 75% completeness for objects brighter than 18 mag, mag, mag, respectively (see also Fremling et al., 2020). The vast majority of our targets (79%) are covered by the BTS program. The remaining targets are usually fainter than 19 mag (see Smith et al.).
@@
This DR2 consists of SNe Ia observed by ZTF between March 2018 and December 2020, with the date cut-off based on the discovery date of the SN. The composition of this sample is detailed in Table 1 and reviewed below.
@@
To build this dataset, we start from the list of any target that has been flagged, at some point, as a “SN Ia” either from our internal databases (Kasliwal et al., 2019; Duev et al., 2019; van der Walt et al.,, 2019; Coughlin et al., 2023), or through the Transient Name Server222www.wis-tns.org. This ‘master list’ contains 3795 targets and of these, 17 have been removed since they have no ZTF light curves but happened to be in our databases.
@@
Since we aim at providing spectroscopically confirmed SNe Ia, we request to be able to release a spectrum that lead to a classification for each target. Therefore, 110 objects flagged as “SN Ia” without a spectrum, or based on non publicly available spectra have been discarded from this release. This corresponds to 2.9% of the initial targets with ZTF light curves. Of the remaining 3668 SNe with at least one spectrum, there were 40 for which a secure Ia classification was not possible. We are thus releasing data for 3628 spectroscopically confirmed SNe Ia.
@@
Following a careful study of the light-curve fit residuals, we concluded in Rigault et al. (2024,b ZTFDR2) that the rest-frame phase range, is sufficiently trained to derive reasonable light-curve model fits. Therefore, considering this phase-range only, we defined targets with “good sampling” as those having, at least: detections at seven phases, two of which are before and two of which are after peak luminosity (), and detections in two photometric bands. Here, we refer to as “phases”, rest-frame phase with respect to estimated maximum light (, see Section 4.7), while same-night same-band detections are ignored, such that, e.g., four detections on the same night in only account for 1 “phase”. These strict phase coverage criteria reduces the number of target by 18.4%, leaving 2960 SNe Ia.
@@
The norm of the SN Cosmology fields usually relaxes the “two pre-max detection” criteria, to “a detection prior to days” (e.g., Betoule et al., 2014; Scolnic et al., 2022). Doing so would leave 3244 objects. However, good sampling pre- and post-max is preferred to ensure a correct estimation of the light curve parameters. We illustrate in Fig. 1 an example of a SN Ia with among the worst light-curve sampling (seven phases), along with SNe Ia with average (40 phases) and best sampled (130 phases) light curves (see light-curve coverage statistics in Smith et al.).
@@
We use the SALT2 light curve fitter (Guy et al., 2010; Betoule et al., 2014; Taylor et al., 2021) to fit our light curves. To ensure the resulting parameters are reasonable for use in cosmological measurements, we further implement the following criteria (see Table 1): ; ; ; ; ; quality of fit (“fitprob”) ; with the measured errors on the SALT2 parameters and with fitprob derived from the best SALT fit (see e.g. Scolnic et al., 2018). This leaves 2667 SNe Ia passing all our cosmology-ready quality cuts. Using the aforementioned looser light-curve coverage criteria that has been used in previous cosmology studies, we would have 2849 targets instead.
@@
Of these 2667 SNe Ia, 994 have a redshift lower than up to where our sample of non-peculiar SNe Ia is free from non-random selection functions, aka a “volume-limited sample” (see Amenouche et al. and detailed study of this sample in Ginolin et al. 2024,a ZTFDR2 and Ginolin et al. 2024,b ZTFDR2): up to the observed distribution of any parameter (e.g., stretch, color, peak-magnitude) is representative of the parameter’s underlying distribution. Hence, non-peculiar ZTF DR2 SNe Ia up to are free from Malmquist bias, and the parameter distributions and their correlations should be representative of that given by nature. See dedicated discussion on Section 4.6.
@@
Careful sub-classification indicates that 2625 of these 2667 SNe Ia are suitable for SN cosmology, i.e. non-peculiar objects. This involves removing peculiar sub-classes like “91bg-like” and “Ia-csm” objects, but keeping sub-classes generally used in cosmological studies, like “91T-like” or “99aa-like” SNe Ia. For further details of these sub-classifications and how they were determined see Dimitriadis et al. and Burgaz et al. (2024,a ZTFDR2).

### v2

@@
We present the first homogeneous release of several thousand spectroscopically classified type Ia supernovae (SNe Ia) with spectroscopic redshifts. This release, named “DR2,” contains 3628 nearby () SNe Ia discovered, followed, and classified by the Zwicky Transient Facility survey between March 2018 and December 2020. Of these, 3000 have good-to-excellent sampling and 2667 pass standard cosmology light curve quality cuts. This release is thus the largest SN Ia release to date, increasing by an order of magnitude the number of well-characterized low-redshift objects. With DR2, we also provide a volume-limited () sample of nearly a thousand SNe Ia. With such a large, homogeneous, and well-controlled dataset, we are studying key current questions on SN cosmology, such as the linearity SNe Ia standardization, the SN and host dependencies, the diversity of the SN Ia population, and the accuracy of current light curve modeling. These, and more, are studied in detail in a series of articles associated with this release. Alongside the SN Ia parameters, we publish our forced-photometry gri-band light curves, 5138 spectra, local and global host properties, observing logs, and a Python tool to facilitate the use and access of these data. The photometric accuracy of DR2 is not yet suited for cosmological parameter inference, which will follow as the “DR2.5” release. We nonetheless demonstrate that our Hubble diagram of several thousands of SNe Ia has a typical 0.15 mag scatter.
@@
For the last thirty years, type Ia supernovae (SNe Ia) have played a central role in building the current standard model of cosmology. In the late 1990s, SNe Ia led to the discovery of the acceleration of the expansion rate of the Universe, the cause of which was dubbed dark energy (Riess et al., 1998; Perlmutter et al., 1999) (see Goobar & Leibundgut (2011) for a review). The existence of dark energy has since been confirmed with high precision by various other cosmological probes, such as the anisotropies in the cosmic microwave background (Spergel et al., 2003; Planck Collab., 2020) and from baryon acoustic oscillations (e.g., Eisenstein et al., 2005; BOSS Collab., 2017). The two decades that followed were those of the SN cosmology field maturation. Many low- () and high-redshift () surveys have enabled us to gather, altogether, SNe Ia. Joining these datasets, the dark energy equation of state parameter, , has been shown to be in good agreement with , which is expected if the acceleration of the Universe’s expansion is due to a simple cosmological constant, , in Einstein’s general relativity equations (Astier et al., 2006; Betoule et al., 2014; Scolnic et al., 2018; Brout et al., 2022). This gain in statistics allows us to test the use of SNe Ia as accurate standard candles of cosmology, since, despite the great success of SN cosmology, their underlying astrophysics are still largely unknown.
@@
It is generally accepted that an SN Ia is the transient event resulting from the thermonuclear explosion of a carbon-oxygen white dwarf (Whelan & Iben, 1973; Iben & Tutukov, 1984; Nugent et al., 2011), triggered by accreting material from a companion star in a binary system (Liu et al., 2023). However, the nature of the companion star (another white dwarf, main-sequence star, etc.) and the explosion mechanism (at or below the Chandrasekhar mass limit) is still unclear and no single picture has emerged (Maoz et al., 2014). In recent years, sample studies of SNe Ia have been used, as they provide a population-wide approach to probe various aspects of SNe Ia physics (e.g. Maguire et al., 2014; Silverman et al., 2015; Papadogiannakis et al., 2019; Tucker et al., 2020; Desai et al., 2024). However, these analyses are so far limited by small-number statistics and/or survey design (e.g., coverage, cadence, and depth). The astrophysical origin of SNe Ia, and their homogeneity, are consequently still highly debated (Jha et al., 2019). Additional data, notably spectroscopic and close to the explosion epoch, would be valuable to discriminate competitive models (e.g., Deckers et al., 2022, and references there in). Nonetheless, understanding the SN Ia mechanism is not an actual requirement for SN cosmology, as long as the astrophysical dependencies can be controlled below the statistical uncertainties. But, without a better grasp on the actual SN Ia astrophysics, such an assertion is difficult.
@@
Since the mid-90s, the success of SN cosmology has emerged from our ability to standardize their brightness from a natural scatter of down to by exploiting two empirical linear relations that correlate the SN Ia’s stretch and color, derived from their light curve, with their absolute peak brightness (Riess et al., 1996; Tripp, 1998; Guy et al., 2010). A scatter of — that is, 7% precision in distance — makes it one of the most precise standard candles in astrophysics. However, only half of this variance can be explained by measurement or light curve modeling errors. Thus, to further control any unexplained astrophysical dependency in SN Ia distances, the correlations between the SN properties and those of their host environments have been extensively studied over the past decade.
@@
Early on, it was suggested that two populations of SNe Ia must coexist to explain the relative rates between progenitors (e.g., Mannucci et al., 2005; Sullivan et al., 2006; Smith et al., 2012): one population follows recent star formation (called “prompt”) with the other related to stellar mass (i.e., an old population of stars, called “delayed”). More delicate for cosmology, the stretch- and color-standardized SN Ia magnitude was then shown to depend on its environment, such that those from low-mass hosts are fainter then those from high-mass galaxies (e.g., Sullivan et al., 2010; Childress et al., 2013; Roman et al., 2018). This so called mass-step has since been used as a third standardization parameter in all cosmological analysis (e.g., Betoule et al., 2014; Scolnic et al., 2018; Brout et al., 2022; Riess et al., 2022). Yet, the origin of this bias is still highly debated. An inaccurate correction of this effect may lead to significant bias in the derivation of cosmological parameters (Rigault et al., 2015, 2020; Smith et al., 2020). Currently, the most discussed models are: differences between the host interstellar dust properties (Brout & Scolnic, 2021; Popovic et al., 2023), or the progenitor age, in the context of the prompt versus delayed model (Rigault et al., 2013, 2020; Nicolas et al., 2021; Briday et al., 2022). Both may actually be true (e.g., Kelsey et al., 2021; Wiseman et al., 2022).
@@
Today, state-of-the-art compilations of samples nearly reach 2000 SNe Ia and are limited by systematic uncertainties (Brout et al., 2022; Vincenzi et al., 2024). Alongside the astrophysical biases, the dominating source of known uncertainties are calibration issues largely due to the requirement to compile SN Ia samples from multiple surveys spanning various redshift ranges (e.g., Betoule et al., 2014; Brout et al., 2022; Vincenzi et al., 2024). This issue is particularly critical at a low-redshift of , at which dozen of samples are merged, the largest of which contains fewer than 200 targets (e.g., Brout et al., 2022).
@@
After two decades, the standard model of cosmology is starting to see the first hints of inconsistencies and, again, SNe Ia are playing a key role. First and foremost, the direct measurement of the Hubble constant, , derived by anchoring the absolute SN Ia standardized luminosity with Cepheids, is found to be higher than theoretical expectations anchored by cosmic microwave background data (e.g., Riess et al., 2016, 2022; Planck Collab., 2020). Most recently, two independent SN Ia compilations concluded that might differ from at the 2 to level (Rubin et al., 2023; DES Collab., 2024). These results, strengthened by the recent DESI Year-1 release (DESI Collab., 2024), could be the sign of new fundamental physics, or hints of unknown systematic biases in the SN Ia distances.
@@
In that context, the Zwicky Transient Facility (ZTF; Bellm et al., 2019; Graham et al., 2019) has been collecting thousands of nearby SNe a year in the northern sky since it started science operations in March 2018. As of mid-2024, ZTF-discovered SNe have been spectroscopically classified. In this paper, we present an overview of the second data release associated with the “Type Ia Supernovae & Cosmology” science working group, aka “DR2.” This dataset contains 3628 type Ia supernovae (SNe Ia) detected, followed, and classified by the ZTF survey and discovered before the end of December 2020.
@@
This DR2 follows Dhawan et al. (2022), in which we illustrate the characteristics of the survey using the first months of operations. Along with this release overview, 20 companion papers that study the dataset, providing key insights on nearby SN Ia physics and their use as cosmological probes, have been released. However, we warn the user that the current light curve photometry calibration does not yet reach the accuracy to unlock cosmological analysis. Any other scientific study is welcomed. Ongoing work is being finished on the survey calibration that will lead to a cosmological parameter inference as the next release: “DR2.5.”
@@
With 2667 well-sampled nearby spectroscopic SNe Ia passing the usual light curve quality cuts, this release is the largest SN Ia release to date, over any redshift range. It increases by an order of magnitude the current state-of-the-art low-redshift sample compilation of 192 targets used in DES Collab. (2024) to anchor their Hubble diagram, and by five the number of SNe Ia ever used for cosmology (Scolnic et al., 2022).
@@
We summarize in Sect. 2 the ZTF survey operation during the period covered by this release. In Sect. 3, we introduce the DR2 sample, and we briefly review the release data in Sect. 4. We review the 20 companion papers in Sect. 5, while the content and access of the release are summarized in Sect. 6. We conclude in Sect. 7.
@@ 2 Summary of the ZTF operation
The ZTF survey employs the ZTF camera mounted on the P48 Schmidt telescope at Mount Palomar Observatory. As is detailed in Dekany et al. (2020), this 576 megapixel camera is made of sixteen 6144×6160 e2v CCD231-C6 charge-coupled devices (CCD) and equipped with three filters: ztf:g (), ztf:r (), and ztf:i (). Each CCD is subdivided into four read-out channels (called quadrants) that have square 1.01 arcsec pixels, selected to match the typical site image quality of 2 arcsec full width at half maximum (see Smith et al. in prep.). The camera has a 47 deg2 field of view with an 86.7% filling factor. It reaches a typical 20.5 mag -limit depth in 30 s, and, with a slightly less than 9 s read-out overhead made while slewing, ZTF is able to cover, at that depth, 3750 deg2/hour. Between March 2018 and December 2020, ZTF acquired 480,572 exposures, with a typical same-filter cadence of three days, thanks to the National Science Foundation (NSF) -funded Mid-Scale Innovations Program (MSIP) public survey (Bellm et al., 2019). During the DR2 operations, the ZTF collaboration also operated a number of additional surveys, including an extragalactic high-cadence survey, with six visits a night in the g and r bands, a larger-area ( deg2) survey that acquired same-night exposures, mostly in and , and an i-band survey with a four-day cadence (Bellm et al., 2019).
ZTF also has a low-resolution integral field spectrograph (SEDm, Blagorodnova et al., 2018; Rigault et al., 2019; Kim et al., 2022; Lezmy et al., 2022) dedicated to spectroscopically classifying transients detected by the photometric survey. The Bright Transient Survey (BTS) ZTF cross working group is designed to use the SEDm, and other spectrographs, to construct a magnitude-limited spectroscopically classified SN sample. As is detailed in Perley et al. (2020), they reach a 97%, 93%, and 75% completeness for objects brighter than 18 mag, mag, and mag, respectively (see also Fremling et al., 2020). The vast majority of our targets (79%) are covered by the BTS program. The remaining targets are usually fainter than 19 mag (see Smith et al. in prep.).
@@
This DR2 consists of SNe Ia observed by ZTF between March 2018 and December 2020, with the date cutoff based on the discovery date of the SN. The composition of this sample is detailed in Table 1 and reviewed below.
@@
To build this dataset, we started from the list of any target that has been flagged, at some point, as an “SN Ia” either from our internal databases (Kasliwal et al., 2019; Duev et al., 2019; van der Walt et al.,, 2019; Coughlin et al., 2023), or through the Transient Name Server.222www.wis-tns.org This “master list” contains 3795 targets and, of these, 17 have been removed, since they have no ZTF light curves but happened to be in our databases.
@@
Since we aim to provide spectroscopically confirmed SNe Ia, we request to be able to release a spectrum that leads to a classification for each target. Therefore, 110 objects flagged as “SN Ia” without a spectrum, or based on non-publicly available spectra, have been discarded from this release. This corresponds to 2.9% of the initial targets with ZTF light curves. Of the remaining 3668 SNe with at least one spectrum, there were 40 for which a secure Ia classification was not possible. We are thus releasing data for 3628 spectroscopically confirmed SNe Ia.
@@
Following a careful study of the light curve fit residuals, we concluded in Rigault et al. (2024,b) that the rest-frame phase range, , is sufficiently trained to derive reasonable light curve model fits. Therefore, considering this phase-range only, we defined targets with “good sampling” as those that have, at least: detections at seven phases, two of which are before and two of which are after peak luminosity (), and detections in two photometric bands. Here, we refer to as “phases” the rest-frame phase with respect to the estimated maximum light (; see Sect. 4.7), while same-night same-band detections are ignored, such that, for example, four detections on the same night in only account for 1 phase. These strict phase coverage criteria reduce the number of targets by 18.4%, leaving 2960 SNe Ia.
@@
The norm of the SN cosmology fields usually relaxes the “two pre-max detection” criteria to “a detection prior to days” (e.g., Betoule et al., 2014; Scolnic et al., 2022). Doing so would leave 3244 objects. However, good sampling pre- and post-maximum is preferred to ensure the correct estimation of the light curve parameters. We illustrate in Fig. 1 an example of an SN Ia with among the worst light curve sampling (seven phases), along with SNe Ia with average (40 phases) and best sampled (130 phases) light curves (see light curve coverage statistics in Smith et al. in prep).
@@
We used the SALT2 light curve fitter (Guy et al., 2010; Betoule et al., 2014; Taylor et al., 2021) to fit our light curves. To ensure the resulting parameters are reasonable for use in cosmological measurements, we further implemented the following criteria (see Table 1): ; ; ; ; ; a quality of fit (“fitprob”) ; with the measured errors on the SALT2 parameters and with fitprob derived from the best SALT fit (see e.g. Scolnic et al., 2018). This left 2667 SNe Ia that pass all our cosmology-ready quality cuts. Using the aforementioned looser light curve coverage criteria that has been used in previous cosmology studies, we would have 2849 targets instead.
@@
Of these 2667 SNe Ia, 994 have a redshift lower than , up to which our sample of non-peculiar SNe Ia is free from nonrandom selection functions, aka a “volume-limited sample” (see Amenouche et al. 2024 and detailed study of this sample in Ginolin et al. 2024,a and Ginolin et al. 2024,b): up to the observed distribution of any parameter (e.g., stretch, color, peak-magnitude) is representative of the parameter’s underlying distribution. Hence, non-peculiar ZTF DR2 SNe Ia up to are free from Malmquist bias, and the parameter distributions and their correlations should be representative of that given by nature (see dedicated discussion on Sect. 4.6).
@@
Careful subclassification indicates that 2625 of these 2667 SNe Ia are suitable for SN cosmology; that is, non-peculiar objects. This involves removing peculiar subclasses like “91bg-like” and “Ia-csm” objects, but keeping subclasses generally used in cosmological studies, like “91T-like” or “99aa-like” SNe Ia (for further details of these subclassifications and how they were determined, see Dimitriadis et al. (2024) and Burgaz et al. (2024,b)).